Circular No. 8812

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF 2003 AZ_{84}, (50000), (55637), AND (90482) M. E. Brown and T.-A. Suer, California Institute of Technology, re-

port the discoveries of four satellites of large transneptunian objects from observations with the Hubble Space Telescope High Resolution Camera (+ F606W filter). Observations of (55637) 2002 UX_{25} on 2005 Aug. 26.64 UT detected a satellite that was 2.5 ± 0.2 mag fainter than the primary at a separation of 0′′.164 ± 0′′.003 in p.a. 153^{o} ± 2^{o}. Observations of (90482) Orcus on 2005 Nov. 13.13 detected a satellite that was 2.7 ± 1 magnitude fainter than the primary at a separation of 0′′.25 ± 0′′.01 in p.a. 128^{o} ± 1^{o}. Observations of 2003 AZ_{84} on 2005 Dec. 2.99 detected a satellite that was 5.0 ± 0.3 mag fainter than the primary at a separation of 0′′.22 ± 0′′.01 in p.a. 321^{o} ± 2^{o}. Observations of (50000) Quaoar on 2006 Feb. 14.90 detected a satellite that was 5.6 ± 0.2 mag fainter than the primary at a separation of 0′′.35 ± 0′′.01 in p.a. 110^{o} ± 1^{o}.

V1281 SCORPII H. Naito and S. Narusawa, Nishi-Harima Astronomical Observatory;

and H. Yamaoka, Kyushu University, report that a low-resolution spectrogram (range 410–670 nm; resolution 1600 at Hα) of the possible nova reported on IAUC 8810 was obtained on Feb. 21.84 UT with the 2.0-m NAYUTA telescope (+ MALLS). The spectrum shows a broad Hα line with a P-Cyg profile (the FWHM of the emission component is ∼ 1800 km/s), along with many other weaker lines, indicating that the star is indeed a classical nova. A narrow Na I D absorption feature suggests the existence of considerable interstellar reddening. Yamaoka adds that Y. Fujita (Kuma-Kogen, Ehime, Japan) made a prediscovery CCD image of the area of the nova with a Nikon D50 camera (+ 105-mm f/2 camera lens) on Feb. 18.850, and the nova was not visible to limiting mag 11.6.

N. N. Samus, Institute of Astronomy, Russian Academy of Sciences, reports that this nova has been given the designation V1281 Sco.

V1280 SCORPII C. Buil, Castanet, France, reports that an optical spectrogram of

V1280 Sco obtained on Feb. 20.20 UT using a 0.28-m Schmidt-Cassegrain telescope (+ Lhires spectrograph) shows an Hα P-Cyg profile and a FWHM of 730 km/s; the intensity of the Hα peak intensity is 2.3 times the local 670-nm continuum. CCD magnitudes by J. D. West, Mulvane, KS: Feb. 21.467 UT, H = 2.87 ± 0.06; 21.483, V = 4.73 ± 0.01.

2007 February 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green